# RoSIS: Robust Framework for Text-Promptable Surgical Instrument Segmentation Using Vision-Language Fusion

### v1

@@ Abstract
Surgical instrument segmentation (SIS) is an essential task in computer-assisted surgeries, with deep learning-based research improving accuracy in complex environments. Recently, text-promptable segmentation methods have been introduced to generate masks based on text prompts describing target objects. However, these methods assume that the object described by a given text prompt exists in the scene. This results in mask generation whenever a related text prompt is provided, even if the object is absent from the image. Existing methods handle this by using prompts only for objects known to be present in the image, which introduces inaccessible information in a vision-based method setting and results in unfair comparisons. For fair comparison, we redefine existing text-promptable SIS settings to robust conditions, called Robust text-promptable SIS (R-SIS), designed to forward prompts of all classes and determine the existence of an object from a given text prompt for the fair comparison. Furthermore, we propose a novel framework, Robust Surgical Instrument Segmentation (RoSIS), which combines visual and language features for promptable segmentation in the R-SIS setting. RoSIS employs an encoder-decoder architecture with a Multi-Modal Fusion Block (MMFB) and a Selective Gate Block (SGB) to achieve balanced integration of vision and language features. Additionally, we introduce an iterative inference strategy that refines segmentation masks in two steps: an initial pass using name-based prompts, followed by a refinement step using location prompts. Experiments on various datasets and settings demonstrate that RoSIS outperforms existing vision-based and promptable methods under robust conditions.
{IEEEkeywords}
Surgical instrument segmentation, Robust text-promptable surgical instrument segmentation, Referring image segmentation, Instance segmentation, Computer vision
@@ 1 Introduction
[…]
As minimally invasive surgeries [1, 2] have been increasingly adopted due to their various advantages, surgeons now rely on camera-captured views to perform procedures, as they cannot directly see the surgical site. Surgical Instrument Segmentation (SIS) [3] has gained attention as a fundamental technology required to assist surgeons in this surgical vision environment. Various vision-based deep learning methods [4, 5, 6, 7, 8, 9] have achieved impressive performance. However, due to the characteristics of SIS, these methods often rely on datasets of limited scale, resulting in insufficient performance, and they lack the capability in terms of human-robot interaction for robot-assisted surgeries. Recently, language and vision-language model-based methods [10, 11] have been developed to overcome these challenges. They introduced a new approach to SIS called text-promptable surgical instrument segmentation [10] based on Referring Image Segmentation (RIS) [12, 13].
RIS generates an object mask based on a given text description. However, RIS methods typically assume that the object described by the text exists within the image. When an incorrect or unrelated text is provided, the model produces a mismatched mask. To address this issue, Generalized Referring Expression Segmentation (GRES) [14] and Robust Referring Image Segmentation (R-RIS) [15] have been proposed. These methods first classify whether the object described by the text actually exists in the image. If the object is present, they proceed to predict the mask.
A similar issue arises in RIS-based text-promptable SIS methods. These methods define a text sentence by utilizing the instrument’s name along with a detailed description generated from the large language model such as GPT-4 [16]. They then predict a mask that corresponds to the given instrument prompt. However, even when a text prompt for an instrument that is not present in the image is provided, the model still produces an incorrect mask. Existing methods handle this by first identifying the instruments in the image and then providing prompts only for those that are detected. However, it utilizes inaccessible information (the existence of an object) in the evaluation phase, causing an unfair comparison with vision-based models.
@@
Figure 1: Comparison of surgical instrument segmentation method: (1) Vision-based segmentation model only uses the image and generates all classes of segmentation maps; (2) Previous promptable segmentation model uses “not accessible information” to identify categories present in the image and generates segmentation maps only for those categories. This means they are led to settings without robust conditions; (3) Our robust promptable segmentation model uses prompts for all classes, generating segmentation maps for all categories under robust conditions.
Figure 2: Performance comparison on the EndoVis2018 dataset under with and without robust conditions. Existing models exhibit significant performance degradation, with dramatic drops in both Challenge IoU (Ch IoU) and ISI IoU. In contrast, our robust promptable segmentation model demonstrates minimal performance degradation across both metrics and outperforms the vision-based SOTA method (ISINet).
Accordingly, inspired by the R-RIS [15] approach, we redefine the text-promptable SIS under robust conditions as called Robust text-promptable Surgical Instrument Segmentation (R-SIS). As illustrated in Fig. 1, previous promptable SIS models [10, 17, 11] use prompts only for categories present within the image, which is inaccessible information in the vision-based models. They generate segmentation maps for a subset of classes because they only use prompts for objects in the image. Unlike previous approaches, R-SIS ensures a fair comparison by incorporating text prompts for all categories within the dataset, regardless of their presence in the image. To achieve this robustness, R-SIS employs negative prompts for non-existing objects and positive prompts for existing objects in training, enabling the model to classify each object’s presence or absence effectively. This strategy avoids false positives and improves accuracy in diverse and realistic surgical scenarios. To support the development and evaluation of R-SIS, we build the dataset for R-SIS using the EndoVis2017 [18] and EndoVis2018 [19] datasets. This adapted dataset includes comprehensive prompts for R-SIS, facilitating more reliable training and evaluation under robust conditions.
Previous promptable SIS methods assume the presence of an object matching the given text, which results in increased false positives in the R-SIS environment. As shown in Fig. 2, previous methods suffer from significant performance degradation when evaluated under robust conditions. There are even cases where the performance is lower than that of vision-based methods that do not use language information. To address this limitation, we propose a novel framework designed to perform well in robust conditions. Our model, named Robust Surgical Instrument Segmentation (RoSIS), includes two key modules: the Multi-Modal Fusion Block (MMFB) and the Selective Gate Block (SGB). The MMFB effectively integrates visual and language features to enhance contextual understanding with language tokens, while the SGB dynamically adjusts the balance of visual and fused features. We also propose the iterative inference to generate more accurate segmentation maps at test time. Iterative inference produces a refined map through two iterations: the first iteration uses prompts based on the object name and a generative model, while the second iteration incorporates location prompts generated using the initial map. This architecture and strategy enable RoSIS to achieve stable and accurate segmentation across a broad range of surgical conditions.
Our model demonstrates competitive performance on the EndoVis2017 and EndoVis2018, achieving powerful results to vision-based models under robust evaluation conditions. To ensure this robustness, we carefully revisit our experimental process, allowing for precise benchmarking. Additionally, we conduct comparisons with existing promptable methods under the R-RIS setting to further validate our approach. This thorough evaluation shows the robustness and effectiveness of our model, highlighting its ability to maintain high accuracy and reliability, even when directly compared with vision-based and promptable segmentation methods.
@@ 2.1 Surgical Instrument Segmentation
Surgical instrument segmentation (SIS) involves predicting the pixel-wise regions of instruments in surgical images. Early research [4, 7, 20, 9, 8] in this field applied image-based segmentation methods to address SIS. TernausNet [4] marked the beginning of deep learning-based approaches by introducing a modified U-Net [21] architecture based semantic segmentation framework with pre-trained VGG11/VGG16 encoders and skip connections for precise location. ISINet [7], employing a Mask R-CNN [22] architecture, incorporates a novel temporal consistency module to track and correct instrument predictions across video frames. TraSeTR [20] utilizes a Track-to-Segment Transformer architecture, combining query embeddings with contrastive learning to ensure robust tracking and segmentation across video frames. MATIS [9], a video-based SIS method, leverages a two-stage structure with Mask2Former [23] to generate and classify region proposals using masked and deformable attention and video transformers to maintain temporal consistency. S3Net [8] builds on Mask R-CNN [22] with a three-stage architecture, adding a Multi-Scale Mask-Attended classifier to refine class predictions based on mask-attended features, improving instance segmentation of surgical instruments. Recently, with the development of segmentation techniques that utilize information beyond vision, promptable segmentation models that use external prompts, such as text or prototypes, have been proposed for SIS.
@@
SurgicalSAM [17] proposes an efficiently fine-tuning method for Segment Anything Model (SAM) [24]. SAM is a segmentation model using prompts (points, bounding boxes, or text inputs) from prompt encoder. Since SAM was originally designed for natural domain image segmentation, SurgicalSAM adapts SAM for the surgical domain, making it suitable for SIS. Additionally, SurgicalSAM introduces a prototype-based class prompt encoder that generates prompt embeddings for each class, which are then used as SAM prompts. Later, various text-promptable SIS methods were proposed, leveraging vision-language models (VLM) to use text prompts for segmentation tasks. SP-SAM [11] uses SAM [24] as its baseline and incorporates both category-level and part-level text prompts as prompt embeddings. TP-SIS [10], a method based on CLIP [25], utilizes text prompts describing instruments, which are categorized into name-based prompts, CLIP’s prompt templates, and prompts generated from GPT-4 [16]. Although these methods have been proposed for text-promptable SIS, they are trained and inferred with the assumption that the object described by the text expression exists within the image. As a result, their performance degrades significantly when a text prompt refers to an object not present in the image. In this paper, we address this issue by adopting a robust referring image segmentation [15] approach for text-promptable SIS. We first determine the existence of the object to reduce false positives and handle negative expressions. Furthermore, we resolve the unfair comparisons that occur in existing text-promptable segmentation models, ensuring a fair comparison with vision-based segmentation models.
@@ 2.2 Referring Image Segmentation
Referring Image Segmentation (RIS) aims to segment the object corresponding to a given text expression. MAttNet [13] designs a language attention network using word embeddings and an LSTM, then compares these with object proposals to generate a mask that matches the expression. VLT [26] and EFN [27] propose a Bi-GRU language model and a ResNet-based image backbone to generate features fused in the decoder through multi-modal information integration. With the development of Vision Transformers (ViT) [28], various methods are applied ViT to RIS tasks. LAVT [12] introduced an early fusion method that combines text features generated by BERT [29] with visual features from ViT through pixel-word attention, achieving significant performance improvements.
However, previous works assumed that each expression must always match an object, which led to the crucial issue of generating incorrect masks for expressions that do not correspond to any object in the image. ReLA [14] overcomes this by proposing Generalized Referring Expression Segmentation, incorporating empty-target (no target for given text) and multi-target annotations (multiple targets for given text). RefSegformer [15] focuses on the empty-target scenario, proposing a robust referring image segmentation approach. The empty-target problem also arises in existing promptable surgical instrument segmentation methods, and we aim to address this issue in our work.
Figure 3: RoSIS model architecture with (a) overall structure, (b) Multi-Modal Fusion Block (MMFB), and (c) Selective Gate Block (SGB).
@@ 3 Method
This section presents our approach to robust text-promptable surgical instrument segmentation (R-SIS). Section 3.1 formulates the problem for R-SIS, identifying key challenges in promptable segmentation for surgical instruments. Section 3.2 defines the construction of text prompts used for training, detailing the use of descriptive prompts to enhance model understanding. Section 3.3 describes the model architecture, including the encoder-decoder structure and multi-modal fusion components. Finally, Section 3.3.2 introduces an iterative inference strategy aimed at refining mask accuracy through repeated inference steps.
@@ 3.1 Problem formulation
Surgical instrument segmentation (SIS) is the task of segmenting surgical tools in an endoscopic surgery image $I\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ denote the height and width of the image, respectively. For text-promptable SIS, we incorporate an image $I$ and a text prompt $T$ containing the names or attributes of the surgical tools, making the segmentation task responsive to specific prompts.
In traditional vision-based SIS, the model takes only image $I$ as input and predicts a mask $M\in\mathbb{R}^{H\times W\times C}$ that includes $C$ classes, each corresponding to a specific surgical instrument. However, in text-promptable SIS, both the image $I$ and a class-specific text prompt $t_{c}$ are provided as input for each class $c$. This enables the model to generate a binary mask $M_{c}\in\mathbb{R}^{H\times W\times 1}$ for each class $c$. For the robust text-promptable surgical instrument segmentation (R-RIS), we also generate an existence probability $p_{c}$, which classifies the class $c$ in the image for a given text prompt $t_{c}$.
@@
To ensure R-SIS, we incorporate existence probabilities, allowing our model to assess the presence of each instrument class first. This setup enables to avoidance of producing unnecessary masks for absent classes, reducing false positives, and enhancing segmentation accuracy under the R-SIS setting. By integrating class existence probabilities, our method is evaluated under robust conditions with vision-based models and demonstrates results under practical conditions.
@@ 3.2 Prompt generation
There are various ways to create text prompts $t_{c}$ for each class $c$. Following the approach in TP-SIS [10], we generate prompts using the class name, such as “The {cls name}” and we utilize GPT-4 [16] to generate detailed prompts describing each class’s visual characteristics. This combination provides both general and specific prompts, enabling the model to recognize instruments based on both name and appearance.
Since surgical instruments often appear in consistent positions on the screen during surgery, we further enhance the prompts by designing location-based descriptions. To do this, we calculate the center of mass for each instrument’s ground truth mask to identify its typical position, creating prompts like “The {cls name} on the {location}.” For location prompts, we categorize positions into four options: “left-top,” “left-bottom,” “right-top,” and “right-bottom.” During model training, we employ all three types of text prompts—class name prompt, GPT-4 visual prompt, and location prompt—for each class $c$, ensuring comprehensive contextual information.
If training only with positive prompts (describing objects present in the image), the model may develop a bias toward the presence of instruments, potentially reducing generalization [15]. To counter this, we also generate negative prompts for classes absent from the image, thus enhancing the classification ability to distinguish between existing and non-existing classes. For absent classes, we create prompts using both the class name and GPT-4 description prompts, as well as a randomly selected location prompt from the four options, to provide a full range of negative examples. We balance the positive and negative prompts corresponding to one image for stable training.
Figure 4: Iterative segmentation process: (Iteration 1) initial segmentation using a general prompt, (Iteration 2) refined segmentation with a location prompt referring to the first predicted map. Then, we combine maps by equation 4 for accurate final output.

### v3

# Rethinking Text-Promptable Surgical Instrument Segmentation with Robust Framework

@@ Abstract
Surgical instrument segmentation is an essential component of computer-assisted and robotic surgery systems. Vision-based segmentation models typically produce outputs limited to a predefined set of instrument categories, which restricts their applicability in interactive systems and robotic task automation. Promptable segmentation methods allow selective predictions based on textual prompts. However, they often rely on the assumption that the instruments present in the scene are already known, and prompts are generated accordingly, limiting their ability to generalize to unseen or dynamically emerging instruments. In practical surgical environments, where instrument existence information is not provided, this assumption does not hold consistently, resulting in false-positive segmentation. To address these limitations, we formulate a new task called Robust text-promptable Surgical Instrument Segmentation (R-SIS). Under this setting, prompts are issued for all candidate categories without access to instrument presence information. R-SIS requires distinguishing which prompts refer to visible instruments and generating masks only when such instruments are explicitly present in the scene. This setting reflects practical conditions where uncertainty in instrument presence is inherent. We evaluate existing segmentation methods under the R-SIS protocol using surgical video datasets and observe substantial false-positive predictions in the absence of ground-truth instruments. These findings demonstrate a mismatch between current evaluation protocols and real-world use cases, and support the need for benchmarks that explicitly account for prompt uncertainty and instrument absence.
keywords:
Surgical instrument segmentation , Text-promptable segmentation , Vision-language models , Robust segmentation , Medical image analysis
[…]
[label1]organization=Center for humanoid research, Korea Institute of Science Technology (KIST),addressline=5, Hwarang-ro 14-gil Seongbuk-gu, city=Seoul, postcode=02792, country=South Korea
@@ 1 Introduction
Minimally invasive surgeries [35, 36] are widely adopted due to their advantages, requiring surgeons to rely on camera-captured views instead of direct visualization. Surgical Instrument Segmentation (SIS) [24] is fundamental to enhancing surgical vision. Vision-based SIS models [2, 31, 32, 3, 6, 5] typically generate predictions for all predefined instrument categories, regardless of whether they are present in the surgical scene. While effective in static settings, this behavior limits integration with interactive systems and automation pipelines, where selective and context-aware segmentation is often required. More recently, promptable segmentation approaches [10, 12] have been introduced, allowing segmentation to be guided by natural language descriptions. Many recent promptable segmentation approaches are inspired by referring image segmentation (RIS) [20, 17], which assumes that a query refers to an object that is present in the image. RIS methods offer selective segmentation over target categories but commonly rely on the assumption that the prompted object is visible in the image. This assumption in RIS also applies to the promptable SIS method.
However, this assumption does not consistently hold in real surgical environments, where we do not know the presence of instruments in the current scene (oracle information). This leads promptable segmentation models to often produce false-positive masks for absent instruments when prompted without presence verification [16]. We need a stricter standard in surgical environments due to the critical risks posed by false-positive masks when instrument presence is uncertain. Moreover, current evaluation protocols for promptable segmentation methods often rely on oracle information that is unavailable in practical scenarios. As illustrated in Fig. 1, these models [10, 11, 12] exclude prompts for absent instruments based on the ground truth, leading to evaluations that do not reflect real-world surgical environments.
To address these limitations, we introduce a new task formulation: Robust text-promptable Surgical Instrument Segmentation (R-SIS). Under this setting, prompts are issued for all candidate categories without assuming knowledge of instrument presence. A model must determine whether the prompted instrument is visually present in the current scene and, only if confirmed, produce an appropriate segmentation mask. This process ensures that segmentation is based on visual features rather than relying on prior assumptions about instrument presence. This setting is also fit for real surgical systems, where the presence of instruments changes frequently in the camera’s view.
@@
Figure 1: Comparison of evaluation process of SIS method: (1) Vision-based segmentation model [2, 31, 32, 3] only uses the image and generates all classes of segmentation maps; (2) Previous promptable segmentation model [10, 11, 12] uses “oracle information” to identify categories present in the image and generates segmentation maps only for those categories. This means they are led to settings without robust conditions; (3) Our R-SIS uses prompts for all classes, generating segmentation maps for all categories under robust conditions.
To support this task, we design a segmentation framework, referred to as Robust Surgical Instrument Segmentation (RoSIS), that integrates visual and textual features, predicts instrument existence, and generates segmentation outputs using multiple prompt types. The framework supports diverse prompt types and introduces iterative refinement to achieve improved segmentation performance under the R-SIS setting, without relying on prior knowledge of instrument presence.
To evaluate segmentation performance under the R-SIS setting, we conduct experiments on two benchmark datasets, EndoVis2017 [25] and EndoVis2018 [26]. Through a series of detailed and diverse experiments, we demonstrate how prompt uncertainty and instrument absence affect segmentation behavior and highlight the importance of realistic evaluation protocols.
In summary, our contributions are three-fold: (1) We define the R-SIS task, which evaluates promptable segmentation under realistic conditions without assuming instrument presence. (2) We present the RoSIS, which is a segmentation framework that supports existence prediction and iterative refinement using diverse prompt types. (3) We establish R-SIS evaluation protocols and benchmarks on two surgical datasets, revealing limitations in existing methods and motivating more robust prompting strategies.
Figure 2: RoSIS model architecture with (a) overall structure, (b) Multi-Modal Fusion Block, and (c) Selective Gate Block.
@@ 2.1 Surgical Instrument Segmentation
SIS involves predicting the pixel-wise regions of instruments in surgical images. Early research [2, 3, 4, 5, 6] applied image-based segmentation methods to SIS. TernausNet [2] introduced a modified U-Net [8] architecture with pre-trained VGG11/VGG16 encoders and skip connections for precise location. ISINet [3], utilizing a Mask R-CNN [7] architecture, incorporates a temporal consistency module to refine instrument predictions across video frames. TraSeTR [4] employs a Track-to-Segment Transformer, combining query embeddings with contrastive learning for robust tracking and segmentation. MATIS [5], a video-based SIS method, uses a two-stage structure with Mask2Former [9] to generate and classify region proposals, incorporating masked and deformable attention and video transformers for temporal consistency. S3Net [6] extends Mask R-CNN [7] with a Multi-Scale Mask-Attended classifier, refining class predictions based on mask-attended features to improve instance segmentation. Recently, with advancements in segmentation techniques utilizing information beyond vision, promptable segmentation models incorporating external prompts, such as text or prototypes, have been proposed for SIS.
@@
SurgicalSAM [11] introduces an efficient fine-tuning method for the Segment Anything Model (SAM) [13], which employs prompts (points, bounding boxes, or text inputs) from a prompt encoder. Since SAM was originally designed for natural domain image segmentation, SurgicalSAM adapts it for the surgical domain. Additionally, it introduces a prototype-based class prompt encoder that generates prompt embeddings for each class as inputs to SAM. Later, various text-promptable SIS methods were proposed, leveraging vision-language models for text-based segmentation. SP-SAM [12] builds upon SAM [13], incorporating both category-level and part-level text prompts. TP-SIS [10], based on CLIP [14], utilizes text prompts describing instruments, categorized into name-based prompts, CLIP’s prompt templates, and prompts generated from GPT-4 [15]. However, these methods assume that the described instrument is always present in the image, leading to incorrect masks when the instrument is absent. To mitigate this, existing approaches first identify the instruments and then generate prompts based on detected instruments. This introduces oracle information during evaluation, leading to unfair comparisons with vision-only models.
We rethink text-promptable SIS by introducing R-SIS, which eliminates this assumption and ensures a fair evaluation. Unlike prior methods that selectively use prompts for detected instruments, R-SIS systematically integrates prompts for all classes, allowing the model to infer instrument presence dynamically. Our approach also addresses the false positives inherent in previous text-promptable SIS models.
@@ 2.2 Referring Image Segmentation
RIS aims to segment objects based on a given text expression. Early works like MAttNet [17] used language attention networks and object proposals to match expressions with objects. Later methods, such as VLT [18] and EFN [19], employed Bi-GRU language models and ResNet-based encoders to fuse multi-modal information. The adoption of Vision Transformers (ViT) [21] has led to further advancements in RIS. LAVT [20] integrates BERT-based text encoding with visual features from ViT through pixel-word attention, significantly improving segmentation accuracy.
However, a key limitation of prior RIS models is the assumption that every text expression corresponds to an object in the image. This assumption leads to incorrect mask generation when no relevant object is present. Generalized Referring Expression Segmentation [23] tackles this issue by introducing empty-target (no target) and multi-target scenarios. RefSegformer [16] further refines this approach by explicitly handling the empty-target problem, ensuring robust segmentation even when the described object is absent. This issue also affects text-promptable SIS methods, where prompts often refer to instruments that do not exist in the image. Our work extends the robust RIS framework to SIS, addressing false positives while ensuring fair comparisons between promptable and vision-based segmentation models.
@@ 3 Method
This section presents our approach to R-SIS and RoSIS. Section 3.1 formulates the problem for R-SIS, identifying the notation and procedure of promptable segmentation for surgical instruments. Section 3.2 defines the construction of text prompts used for training, detailing the use of descriptive prompts to enhance model understanding. Section 3.3 describes the architecture of RoSIS, which is an encoder-decoder structure with multi-modal fusion components and an instrument existence prediction branch. Finally, Section 3.4 introduces an iterative refinement strategy, applied during inference, that is designed to address the R-SIS task by determining the presence of instruments and leveraging diverse prompts to generate refined segmentation masks.
@@ 3.1 Problem formulation
SIS is the task of segmenting surgical tools in an endoscopic surgery image $I\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ denote the height and width of the image, respectively. For text-promptable SIS, we incorporate an image $I$ and a text prompt $T$ containing the names or attributes of the surgical tools, making the segmentation task responsive to specific prompts.
In traditional vision-based SIS, the model takes only image $I$ as input and predicts a mask $M\in\mathbb{R}^{H\times W\times C}$ that includes $C$ classes, each corresponding to a specific surgical instrument. However, in text-promptable SIS, both the image $I$ and a class-specific text prompt $t_{c}$ are provided as input for each class $c$. This enables the model to generate a binary mask $M_{c}\in\mathbb{R}^{H\times W\times 1}$ for each class $c$. For the R-SIS, we also generate an existence probability $p_{c}$, which classifies the class $c$ in the image for a given text prompt $t_{c}$.
@@
To enable R-SIS, we incorporate existence probabilities, allowing our model to assess the presence of each instrument class first. This setup enables to avoidance of producing unnecessary masks for absent classes, reducing false positives, and enhancing segmentation accuracy under the R-SIS setting. By integrating class existence probabilities, our method is evaluated under robust conditions with vision-based models and demonstrates results under practical conditions.
@@ 3.2 Prompt generation
There are various ways to create text prompts $t_{c}$ for each class $c$. Following the approach in [10], we generate prompts using the class name, such as “The {cls name}” and we utilize GPT-4 [15] to create detailed prompts describing each class’s visual characteristics and keep them during the train and test phase. This combination provides both general and specific prompts, enabling the model to recognize instruments based on both name and appearance.
Since surgical instruments often appear in consistent positions on the screen during surgery, we further enhance the prompts by designing location-based descriptions. To do this, we calculate the center of mass for each instrument’s ground truth mask to identify its typical position, creating prompts like “The {cls name} on the {location}.” During training, these location prompts are generated using ground truth annotations to ensure accurate positional information. For location prompts, we categorize positions into four options: “left-top,” “left-bottom,” “right-top,” and “right-bottom.” During model training, we employ all three types of text prompts—class name prompt, GPT-4 visual prompt, and location prompt—for each class $c$, ensuring comprehensive contextual information. During inference, since ground truth is not accessible, we apply an iterative refinement strategy to dynamically estimate instrument locations and construct corresponding location prompts, thereby maximizing their utility under deployment conditions.
If training only with positive prompts (describing instruments present in the image), the model may develop a bias toward the presence of instruments, potentially reducing generalization [16]. To counter this, we also generate negative prompts for classes absent from the image, thus enhancing the classification ability to distinguish between existing and non-existing classes. For absent classes, we create prompts using both the class name and GPT-4 description prompts, as well as a randomly selected location prompt from the four options, to provide a full range of negative examples. We balance the positive and negative prompts corresponding to one image for stable training.